# Vision Transformers Don’t Need Trained Registers

### v4

@@ Appendix A Appendix
We investigate the impact of the number of test-time registers and the initialization strategy in Sections A.1 and A.2. Additional qualitative attention visualizations are provided in Section A.3, followed by further analysis of OpenCLIP register neuron activations in Sections A.5 and A.6. We apply the analysis from Section 3 to DINOv2 in Section A.7. Additional results from the experiments in Section 5 are included in Sections A.8 and A.9. Finally, we provide additional VLM results with test-time registers in Section A.10.
A.1 Ablating number of register tokens
@@
In the main text, we focused on evaluation using a single register. Here, we examine the impact of employing multiple registers. Following the analysis from Sun et al. (2024), we separate the influence of high-norm tokens on the attention output (i.e. after multiplication with value features). The attention output at each token $t$ can be decomposed into two components: value contributions from the register tokens $\mathcal{R}$, and value contributions accumulated over the CLS and patch tokens.
@@
Experimental setup for test-time registers. Given Equation 1, we analyze the contribution of test-time registers to the value updates of all other tokens in the final attention layer of DINOv2 ViT-L/14. We first compute the value update for each token across the ImageNet validation set using a single test-time register. Then, we vary the number of registers and, for each setting, compute the cosine similarity between the resulting value update and the one obtained with a single register, averaged over all tokens. Since all test-time registers are initialized to zero, we break symmetry by randomly assigning which test-time register receives the outlier register neuron activation. Thus, each test-time register holds the activations from a different set of register neurons.
@@
Increasing the number of test-time registers beyond one has a negligible impact. Our results in Table 8 show that increasing the number of registers has a marginal impact on the internal computation of the model. We see the qualitative effect of this on the last layer’s CLS token attention map averaged across all heads in Figure 9. One test-time register removes all the high-norm artifacts from the original model’s attention map. These results align with the finding from Sun et al. (2024) that register tokens act as attention biases. Thus, a single register token can be sufficient to induce the same value update across tokens. We further corroborate this behavior on trained registers next.
Experimental setup for trained registers. We use DINOv2 ViT-L/14 trained with four registers and calculate the value update contributions from each of the four registers to both image patch tokens and the CLS token. To approximate their collective effect with a single token, we construct a synthetic register by assigning, for each embedding dimension, the value (including sign) with the largest absolute magnitude across the four trained registers. We then compute the value update induced by this synthetic register and measure its cosine similarity with the original update from the four-register setup, averaged over all tokens.
@@
Figure 9: One test-time register is sufficient to remove artifacts. We increment the number of test-time registers where each register holds the activations from a different set of register neurons. The last layer’s average CLS token attention map from DINOv2 ViT-L/14 does not significantly change with more test-time registers.
A.2 Evaluating different initial values for test-time registers
@@
Different initializations produce similar results. All three approaches yield similar probing performance as reported in Table 9. We attribute this outcome to the fact that the primary contribution of the register during attention comes from the large activations it holds, while the other values, which are much smaller in magnitude, have little impact on the final result. In Figure 10, we visualize the last layer’s average CLS token attention map from DINOv2 ViT-L/14, and observe that different initialization strategies do not significantly impact the test-time register’s ability to remove artifacts.
A.3 Attention maps after intervening upon register neurons

### v5

@@ Appendix A Appendix
A.1 Proof of Proposition 1 and Corollary 1
Set-Up. We consider the MLP layer from the penultimate Transformer block followed by the full final block (attention + MLP) of a residual Vision Transformer without normalization layers. Let the input sequence to the penultimate MLP be a task-specific [CLS] token and $n$ patch tokens: $x^{(0)}_{cls},x^{(0)}_{1},...x^{(0)}_{n}\in\mathbb{R}^{d}$ The MLPs will consist of two linear layers (no bias) and an activation function $\phi$ (e.g., ReLU or GELU). We denote the weight matrices of the penultimate block’s MLP as $W_{\text{in}}^{(1)}\in\mathbb{R}^{d\times d_{\text{mlp}}},W_{\text{out}}^{(1)}\in\mathbb{R}^{d_{\text{mlp}}\times d}$, and those of the final block’s MLP as $W_{\text{in}}^{(2)}\in\mathbb{R}^{d\times d_{\text{mlp}}},W_{\text{out}}^{(2)}\in\mathbb{R}^{d_{\text{mlp}}\times d}$. There will be one attention head with identity projection matrices: $W_{Q}=W_{K}=W_{V}=W_{O}=I_{d}\in\mathbb{R}^{d\times d}$. Finally, the output task-specific head is a linear projection that uses the [CLS] token’s embedding for prediction: $W_{\text{head}}\in\mathbb{R}^{d\times c}$, where $c$ is the number of output classes (or regression targets).
Our analysis is carried out in the regime where both $W_{\text{in}}^{(2)}$, $W_{\text{head}}$ are low-rank, i.e., $\mathrm{rank}(W_{\text{in}}^{(2)})<\mathrm{min}(d,d_{\text{mlp}})$ and $\mathrm{rank}(W_{\text{head}})<\mathrm{rank}(d,c)$. Hence, both matrices possess non-trivial kernels. This low-rank property is consistent with the effective structure observed in practice (Li et al., 2018; Aghajanyan et al., 2021; Hu et al., 2022).
Proposition 1 (Register neuron induces attention sink and no-op attention)
Let $u_{1}=(W_{\text{in}}^{(1)})_{:,1}$ be a register neuron and $u_{2}=(W_{\text{out}}^{(1)})_{1,:}$ be the corresponding row in the MLP’s second weight matrix, with $\|u_{2}\|\gg\|(W_{\text{out}}^{(1)})_{j,:}\|$ for $j\neq 1$. If both $u_{1},u_{2}\in\ker(W_{\text{in}}^{(2)\top})\cap\ker(W^{\top}_{\text{head}})$ and a token $x_{s}$ is aligned with $u_{1}$, i.e., $x^{(0)}_{s}=\beta u_{1}$ for some $\beta>0$, then $x_{s}$ becomes an attention sink and the attention layer has no contribution to the model’s output.
Proof:
After the first MLP with a residual connection, the updated hidden states for the tokens are:
$x^{(1)}_{i}\;=\;x^{(0)}_{i}\;+\;\phi\big(x^{(0)}_{i}W_{\text{in}}^{(1)}\big)\,W_{\text{out}}^{(1)},\qquad i\in\{cls,1,2,...,n\}$ (2)
This can be interpreted as adding a unit vector $v_{i}\in\mathbb{R}^{d}$ to the tokens modulated by some coefficient $\alpha_{i}$:
$x^{(1)}_{i}\;=\;x^{(0)}_{i}\;+\;\alpha_{i}v_{i},\qquad i\in\{cls,1,2,...,n\}$ (3)
If $x^{(0)}_{s}$ is aligned with the register neuron $u_{1}$ (i.e., $x^{(0)}_{s}=\beta u_{1}$ for some $\beta>0$), its activation $\tilde{\alpha}_{s}=\phi(x_{s}^{(0)\top}u_{1})=\phi(\beta{u_{1}^{\top}}u_{1})$ will positively scale the corresponding row in the second MLP matrix $u_{2}$. As $\|u_{2}\|\gg\|(W_{\text{out}}^{(1)})_{j,:}\|$ for $j\neq 1$, the update is dominated by the direction of $u_{2}$, making it the principal contribution to $x_{s}$ in the residual stream. Thus, $x^{(1)}_{s}\approx x^{(0)}_{s}+\tilde{\alpha}_{s}u_{2}=\beta u_{1}+\tilde{\alpha}_{s}u_{2}$. To make the norm and direction of the update explicit, we rewrite this as $x^{(1)}_{s}=\beta u_{1}+\alpha_{s}v_{2}$, where $v_{2}=\frac{u_{2}}{\|u_{2}\|}$ and $\alpha_{s}=\tilde{\alpha}_{s}\|u_{2}\|.$ As the register neuron activation $\|\tilde{\alpha}_{s}\|\to\infty$, it follows that $\|\alpha_{s}\|,\|\tilde{\alpha}_{s}\|\gg\|\alpha_{i}\|$ for $i\neq s,$ and $\|x^{(1)}_{s}\|\gg\|x^{(1)}_{i}\|$ for $i\neq s$.
Given identity projection matrices during attention (i.e., the key, queries, values are the tokens themselves), the attention update for the [CLS] token is:
$x_{cls}^{(2)}=x_{cls}^{(1)}+\sum_{i\in\{cls,1,\dots,n\}}p_{cls,i}\,x_{i}^{(1)},$ (4)
where $p_{cls,i}$ is the softmax attention weight assigned by the [CLS] token to token $i$:
$p_{cls,i}=\frac{\exp\!\big(x_{cls}^{(1)\top}x_{i}^{(1)}\big)}{\sum_{j\in\{cls,1,\dots,n\}}\exp\!\big(x_{cls}^{(1)\top}x_{j}^{(1)}\big)}.$ (5)
As $\|x^{(1)}_{s}\|\to\infty$, the dot product $x^{(1)\top}_{cls}x^{(1)}_{s}=\|x^{(1)}_{cls}\|\cdot\|x^{(1)}_{s}\|\cdot\cos(\theta)$ grows without bound even for small angle $\theta$ between the two token embeddings, leading the softmax to assign nearly all attention weight to $x^{(1)}_{s}$. Thus,
$x^{(2)}_{cls}\approx x^{(1)}_{cls}+x^{(1)}_{s}=x^{(1)}_{cls}+\beta u_{1}+\alpha_{s}v_{2}$ (6)
is the [CLS] token representation after the attention layer.
Since, $u_{1},v_{2}\in ker(W_{\text{in}}^{(2)\top})$, the [CLS] token representation after the second MLP layer is:
$x^{(3)}_{cls}$ $=x^{(2)}_{cls}+\phi\!\left(x^{(2)}_{cls}W_{\text{in}}^{(2)}\right)W_{\text{out}}^{(2)}$ (7)
$=\big(x^{(1)}_{cls}+\beta u_{1}+\alpha_{s}v_{2}\big)+\phi\big((x^{(1)}_{cls}+\beta u_{1}+\alpha_{s}v_{2})W_{\text{in}}^{(2)}\big)W_{\text{out}}^{(2)}$ (8)
$=x^{(1)}_{cls}+\beta u_{1}+\alpha_{s}v_{2}+\phi\big(x^{(1)}_{cls}W_{\text{in}}^{(2)}\big)W_{\text{out}}^{(2)}.$ (9)
Finally, the output prediction is
$x^{(3)}_{cls}W_{\text{head}}$ $=\big(x^{(1)}_{cls}+\beta u_{1}+\alpha_{s}v_{2}+\phi(x^{(1)}_{cls}W_{\text{in}}^{(2)})W_{\text{out}}^{(2)}\big)W_{\text{head}}$ (10)
$=\big(x^{(1)}_{cls}+\phi(x^{(1)}_{cls}W_{\text{in}}^{(2)})W_{\text{out}}^{(2)}\big)W_{\text{head}},$ (11)
since $u_{1},v_{2}\in\ker(W^{\top}_{\text{head}})$. This is equivalent to the forward pass without the attention layer, thus exhibiting the “no-op” attention behavior.
Remarks. In this proof, we assumed $\|u_{2}\|\gg\|(W_{\text{out}}^{(1)})_{j,:}\|$ for $j\neq 1$. In practice, we do observe that the register neurons’ corresponding rows in the MLP’s second weight matrix have specific high norm dimensions (Figure˜25).
Corollary 1 (Register neuron induces implicit attention bias)
Given the aforementioned set-up, the contribution of register neurons is equivalent to the attention mechanism augmented with explicit bias terms ($\mathbf{k}^{\prime},\mathbf{v}^{\prime}$) as in Sun et al. (2024):
$\text{Attention}(Q,K,V;\mathbf{k}^{\prime},\mathbf{v}^{\prime})=\text{softmax}\left(Q\begin{bmatrix}K^{T}&\mathbf{k}^{\prime}\end{bmatrix}\right)\begin{bmatrix}V\\ \mathbf{v}^{\prime T}\end{bmatrix}$
Proof:
As our set-up uses identity projection matrices during attention (i.e., the key, queries, values are the tokens themselves), the attention mechanism is:
$\text{Attention}(Q,K,V;\mathbf{x}^{\prime},\mathbf{x}^{\prime})=\text{softmax}\!\left(X\begin{bmatrix}X^{\top}&\mathbf{x}^{\prime}\end{bmatrix}\right)\begin{bmatrix}X\\ \mathbf{x}^{\prime\!\top}\end{bmatrix}$. (12)
where $X\in\mathbb{R}^{n\times d}$ is the stacked representations of $n$ tokens. From Equation˜6, the attention update for the [CLS] token can be decomposed into a contribution from the sink token prior to the first MLP and from the register neuron:
$x^{(2)}_{cls}\approx x^{(1)}_{cls}+x^{(1)}_{s}=x^{(1)}_{cls}+\underbrace{\beta u_{1}}_{\text{pre-MLP sink token}}+\underbrace{\alpha_{s}v_{2}}_{\text{register neuron contribution}}$ (13)
Based on the proof for Proposition˜1, as the register neuron activation $\|\tilde{\alpha}_{s}\|\to\infty$, it follows that $\|\alpha_{s}\|,\|\tilde{\alpha}_{s}\|\gg\|\alpha_{i}\|$ for $i\neq s,$ and $\|x^{(1)}_{s}\|\gg\|x^{(1)}_{i}\|$ for $i\neq s$. In other words, as the register neuron activation grows unboundedly, it dominates the sink token representation and the sink token achieves a significantly larger norm than all other tokens. As $\|x^{(1)}_{s}\|\to\infty$, the dot product $x^{(1)\top}_{cls}x^{(1)}_{s}=\|x^{(1)}_{cls}\|\cdot\|x^{(1)}_{s}\|\cdot\cos(\theta)$ grows without bound even for small angle $\theta$ between the two token embeddings, leading the softmax to assign nearly all attention weight to $x^{(1)}_{s}$. Thus, the attention update is
$x^{(2)}_{cls}\approx x^{(1)}_{cls}+x^{(1)}_{s}=x^{(1)}_{cls}+\beta u_{1}+\alpha_{s}v_{2}$ (14)
As the register neuron activation $\|\tilde{\alpha}_{s}\|\to\infty$, it follows that $\|\alpha_{s}\|,\|\tilde{\alpha}_{s}\|\gg\beta$. We take a leading order approximation to simplify Equation˜14 to
$x^{(2)}_{cls}\approx x^{(1)}_{cls}+\alpha_{s}v_{2}$ (15)
A similar update rule holds for any arbitrary token $i$:
$x^{(2)}_{i}\;=\;x^{(1)}_{i}\;+\;\alpha_{s}v_{2},\qquad i\in\{cls,1,2,...,n\}$ (16)
This update can be achieved by plugging in the register neuron contribution $\alpha_{s}v_{2}$ into $\mathbf{x^{\prime}}$ in Equation˜12:
$\text{Attention}(Q,K,V;\alpha_{s}\mathbf{v_{2}},\alpha_{s}\mathbf{v_{2}})=\text{softmax}\!\left(X\begin{bmatrix}X^{\top}&\alpha_{s}\mathbf{v_{2}}\end{bmatrix}\right)\begin{bmatrix}X\\ \alpha_{s}\mathbf{v_{2}}^{\!\top}\end{bmatrix}$. (17)
Using a similar argument from the proof for Proposition˜1, as $\|\alpha_{s}\|\to\infty$, for any token index $i$, the dot product $x^{(1)\top}_{i}(\alpha_{s}v_{2})=\|x^{(1)}_{i}\|\cdot\|\alpha_{s}v_{2}\|\cdot\cos(\theta)$ grows without bound even for small angle $\theta$ between the two vectors, leading the softmax to assign nearly all attention weight to $\alpha_{s}v_{2}$. Thus, the update representation for token $i$ after this augmented attention is:
$x^{(2)}_{i}\;=\;x^{(1)}_{i}\;+\;\alpha_{s}v_{2},\qquad i\in\{cls,1,2,...,n\}$ (18)
which is equivalent to Equation˜16, thus showing that register neurons induce a bias term in the attention mechanism. This suggests that we can directly plug in the contribution of register neurons into the attention mechanism as in eq.˜17, without the need for a test-time register. We investigate this in the next section.
A.2 Removing outliers with attention biases
We propose to create these attention biases in OpenCLIP ViT-B/16 training-free by approximating the effect of a test-time register on the attention computations. To set $\mathbf{k^{\prime}}$ and $\mathbf{v^{\prime}}$ for each attention head, we compute the mean key and value vectors of a test-time register token across 1000 images. At test-time, we zero out the activations of register neurons using the parameters from Section˜3.2.
[…]
Figure 9: Using attention biases produces similarly clean attention maps as using a test-time register. We compute the attention map of the [CLS] token averaged across the attention heads of the last layer on OpenCLIP-B/16, finding that both attention biases and test-time registers remove artifacts.
OpenCLIP Acc.
ViT-B/16 71.3
w/ zeroed register neurons 55.6
w/ attention bias 71.3
A.3 Algorithm overview
We provide a broad overview of Algorithm˜1 and Algorithm˜2 accompanied by illustrations of their execution. To find register neurons with Algorithm˜1, we compute the average activation of each neuron on high-norm patches across an image set and return the top_ktop\_k neurons with the highest values (Figure˜10). Using these discovered register neurons, we implement a test-time register by initializing a dummy token to a vector of zeros which is passed forward along with the regular tokens. According to Algorithm˜2, whenever a register neuron is encountered during the forward pass, the maximum register neuron activation across the tokens is copied into the test-time register and the activations of this register neuron are zeroed-out elsewhere (Figure˜11).
Figure 10: Finding register neurons. Our register neuron discovery algorithm returns the neurons with the highest activations on outlier patch locations.
Figure 11: Test-time register intervention. During the forward pass, the max activation from each layer’s register neuron is shifted to the test-time register. The register neuron activations are then zeroed-out over the other tokens, and the inference computation proceeds.
Algorithm 1 FindRegisterNeurons
1:Input: Image set $\mathcal{I}=\{I_{1},\dots,I_{M}\}$, maximum layer index top_layertop\_layer, number of register neurons to return top_ktop\_k, number of neurons in each layer NN
2:Output: top_ktop\_k register neurons
[…]
4:for all $I_{i}\in\mathcal{I}$ do
5:  $O\leftarrow\mathrm{FindOutliers}(I_{i})$ # get indices of top-norm patches
6:  for $\ell=0$ to top_layertop\_layer do
7:    for $n=0$ to N1N-1 do
8:     $\textit{avg\_act}[\ell,n]\leftarrow\textit{avg\_act}[\ell,n]+\frac{1}{|O|M}\sum_{p\in O}\mathrm{activation}_{\ell,n}(I_{i},p)$ # avg. activation of neuron on outlier patches
9:    end for
10:  end for
11:end for
12:return top_ktop\_k neurons with largest avg_act
Algorithm 2 ShiftRegisterNeurons
1:Input: List of register neuron indices in this layer $\mathcal{R}$, array of $n$ neuron activation maps over $T$ tokens produced by this layer $A\in\mathbb{R}^{T\times N}$
2:Output: updated AA
3:for all $r\in\mathcal{R}$ do
4:  $A[-1,r]\leftarrow\max_{t\in[0,T-1]}A[t,r]$ # set test-time register token (idx = -1) to max register neuron activation value over all tokens
5:  $A[0{:}T-1,r]\leftarrow 0$ # zero-out register neuron activation for all other tokens
6:end for
7:return AA # proceed with forward pass
A.4 Hyperparameter Selection
Algorithm˜1 relies on three hyperparameters: (1) the outlier threshold, (2) the index of the highest layer we search up to when identifying register neurons, and (3) the number of register neurons to return. The first two are straightforward to set using simple heuristics. Specifically, the outlier threshold can be defined using the classical criterion of three standard deviations above the mean patch norm. The second, top_layertop\_layer, is chosen as the layer at which outliers first emerge. Setting the number of neurons to return is slightly more involved, relying on sweeping top_ktop\_k until the outliers are suppressed. However, this is manageable in practice since the number of register neurons is sparse. We test how scalable this approach is in Section˜A.7.
A.5 Ablating number of register tokens
@@
In the main text, we focused on evaluation using a single register. Here, we examine the impact of employing multiple registers. Following the analysis from Sun et al. (2024), we separate the influence of high-norm tokens on the attention output (i.e. after multiplication with value features). The attention output at each token $t$ can be decomposed into two components: value contributions from the register tokens $\mathcal{R}$, and value contributions accumulated over the [CLS] and patch tokens.
@@
Experimental setup for test-time registers. Given Equation˜19, we analyze the contribution of test-time registers to the value updates of all other tokens in the final attention layer of DINOv2 ViT-L/14. We first compute the value update for each token across the ImageNet validation set using a single test-time register. Then, we vary the number of registers and, for each setting, compute the cosine similarity between the resulting value update and the one obtained with a single register, averaged over all tokens. Since all test-time registers are initialized to zero, we break symmetry by randomly assigning which test-time register receives the outlier register neuron activation. Thus, each test-time register holds the activations from a different set of register neurons.
@@
# Test-time Registers NYUd rmse $\downarrow$
0 $0.3876\pm 0.0014$
1 $0.3774\pm 0.0013$
2 $0.3771\pm 0.0014$
3 $0.3785\pm 0.0014$
4 $0.3791\pm 0.0017$
5 $0.3795\pm 0.0016$
Table 9: NYUd RMSE for DINOv2-L/14 with varying numbers of test-time registers. Increasing the number of registers beyond one has minimal gains.
Increasing the number of test-time registers beyond one has a negligible impact. Our results in Table˜9 show that increasing the number of registers has a marginal impact on the internal computation of the model. We see the qualitative effect of this on the last layer’s [CLS] token attention map averaged across all heads in Figure˜12. One test-time register removes all the high-norm artifacts from the original model’s attention map. We then evaluate the impact on downstream performance using a linear probe on the NYUd depth estimation task. Overall, we find that adding more test-time registers beyond one does not significantly change performance. We note that there is a slight degradation in performance beyond three test-time registers, although performance remains higher than without registers. These results align with the finding from Sun et al. (2024) that register tokens act as implicit attention biases. Thus, a single register token can be sufficient to induce the same value update across tokens and clean up internal features. We further corroborate this behavior on trained registers next.
Experimental setup for trained registers. We use DINOv2 ViT-L/14 trained with four registers and calculate the value update contributions from each of the four registers to both image patch tokens and the [CLS] token. To approximate their collective effect with a single token, we construct a synthetic register by assigning, for each embedding dimension, the value (including sign) with the largest absolute magnitude across the four trained registers. We then compute the value update induced by this synthetic register and measure its cosine similarity with the original update from the four-register setup, averaged over all tokens.
@@
Figure 12: One test-time register is sufficient to remove artifacts. We increment the number of test-time registers where each register holds the activations from a different set of register neurons. The last layer’s average [CLS] token attention map from DINOv2 ViT-L/14 does not significantly change with more test-time registers.
A.6 Evaluating different initial values for test-time registers
@@
Different initializations produce similar results. All three approaches yield similar probing performance as reported in Table˜10. We attribute this outcome to the fact that the primary contribution of the register during attention comes from the large activations it holds, while the other values, which are much smaller in magnitude, have little impact on the final result. In Figure˜13, we visualize the last layer’s average [CLS] token attention map from DINOv2 ViT-L/14, and observe that different initialization strategies do not significantly impact the test-time register’s ability to remove artifacts.
A.7 InternViT-6B
To test the scalability of our approach, we apply our analysis from Section˜3 to InternViT-6B, one of the largest publicly available vision transformers (Zhu et al., 2025). We find that outliers form at layer 29 (out of 45). We then tracked the max norms across image patches at the output layer, and found that outlier patch norms reach up to a value of 7000, while the median patch norm value is 332 and the standard deviation is 396. Using Algorithm˜1, we then identified 300 register neurons out of a total of 576,000 candidate neurons responsible for driving outlier formation. When we apply a test-time register with these register neurons, the maximum output patch norm at the output layer only reaches 608 on average, effectively mitigating the outliers.
To assess the downstream impact, we ran unsupervised object discovery using the LOST algorithm (as in Section 5.3) since Darcet et al. (2024) found that the performance on this task correlates with the smoothness of a model’s attention maps. We report the correct localization (corloc) results below with and without a test-time register on InternViT-6B, and observe up to a 13-point correct localization improvement.
Figure 14: Mitigating artifacts in InternViT-6B. Using a test-time register removes artifacts from the [CLS] attention map.
VOC 2007 VOC 2012 COCO 20k
InternViT-6B 30.3 35.4 22.4
w/ test-time register 41.6 48.7 32.7
Table 11: Unsupervised Object Discovery with LOST (Siméoni et al., 2021). Using a test-time register boosts LOST performance on InternViT-6B, one of the largest open-source ViTs.
A.8 Typographic attacks
Darcet et al. (2024) showed that outlier patches retain less local information by demonstrating that their pixel/position information is harder to recover with a linear probe compared to normal patches. However, this analysis leaves open the possibility that local information at outliers might still be redistributed to other patches and indirectly influence the [CLS] token. We demonstrate here that this is not the case and that outlier tokens mask local patch information. To evaluate this masking effect, we use typographic attacks as a test-bed. Azuma & Matsui (2023) showed that CLIP is vulnerable to typographic attacks, where images can be misclassified based on written text in the image rather than the depicted object (Figure˜15). We strategically place high-norm artifacts within an image to mask out adversarial patches while preserving semantic content. This intervention is highly localized and relies on activating only a small fraction of neurons to produce targeted changes in the image representation.
Experimental setup. Following Azuma & Matsui (2023), we use OpenAI CLIP (Radford et al., 2021) and identify register neurons with Algorithm˜1 on this model. Then, we localize the largest region of text in the image with OCR and, using the register neurons, move high-norm outliers onto the patches corresponding to the text location. We evaluate on the RTA-100 dataset (Azuma & Matsui, 2023), which contains approximately 1000 images with text written on top of an object from one of 100 classes. We calculate zero-shot accuracy by comparing CLIP’s image embedding to the “real” and “attack” labels’ text embedding.
Model Attack success %
CLIP 50.5
w/ pixel ablation 7.6
w/ register neuron edit 7.5
w/ random neuron edit 50.5 $\pm$ 0.2
Results. Qualitative results in Figure˜15 show that shifting the activation maps of ten register neurons early in the model’s computation causes outlier patches to later appear corresponding to the text area. Table˜12 presents the attack success rate and find that it significantly drops after our intervention, indicating that our method masks part of the model’s internal representation without harming semantic content. Our results suggest that patch information at outliers is lost from the model computation rather than transferred elsewhere. We also present results of an alternative masking procedure in the input space, setting areas corresponding to text to the mean pixel value of the region. Our intervention matches this method’s performance using only a sparse modification – repurposing CLIP’s internal mechanisms by modifying roughly $0.02\%$ of all neurons, compared to masking ${\sim}10\%$ of the input. This highlights the influence of register neurons in guiding CLIP’s visual representation and output behavior – shifting outliers on top of the text areas is performance-wise equivalent to the text never being present from the start. In contract, editing three random sets of ten neurons has minimal effect (w/random neuron edit). The finding that the register neurons effectively mask patch information explains why mitigating the artifacts can improve dense prediction tasks.
A.9 Attention maps after intervening upon register neurons